# Emergent symmetries in prethermal phases of periodically driven quantum systems

### v1

@@ III.1 Prethermal HSF
where $\hat{A}_{j}=(\hat{n}_{j+2}-\hat{n}_{j-1})$, $\gamma_{1}=V_{1}T/(4\hbar)$ and $\alpha(\gamma_{0})=\gamma_{1}^{-1}\sin\gamma_{1}\exp[i\gamma_{1}\hat{A}_{j}]$. For $\gamma_{1}=n\pi$, or $V_{1}=2n\hbar\omega_{D}$, $\alpha(\gamma_{1})=0$. At these special frequencies, $H_{F}^{(1)}$ represents a model of interacting 1D spinless fermions with constrained hopping; such a model, in equilibrium, leads to strong HSF [104]. However, higher order terms in $H_{F}$ [80],do not respect such constraints. Most importantly, the presence of such constraints lead to two approximate emergent conservations; $[H_{F}^{(1)},n_{d}]=0$ and $[H_{F}^{(1)},n_{d}^{s}]=0$, where $n_{d}=\sum_{j}\hat{n}_{j}\hat{n}_{j+1}$ and $n_{d}^{s}=\sum_{j}(-1)^{j}\hat{n}_{j}\hat{n}_{j+1}$. Note that these conservation laws are approximate as higher order terms in $H_{F}$ [80] do not respect such constraints. These approximate emergent conservation laws are crucial in obtaining a prethermal HSF phase at special frequencies.
@@
The signature for such HSF becomes apparent in several quantities. For example, one expects that entanglement entropy for a generic initial state of an ergodic driven system to eventually saturate to its Page value, $S_{p}$, which can be thought as the average entropy for a generic ergodic many-body system; it is also the typical entropy for large enough systems and depends on the HSD of the system (see Refs. 131, 132 for details). However, at these special frequencies and for high drive amplitude, the half-chain entanglement entropy $S(nT)$, of an initial Fock state after $n$ drive cycles computed using periodic boundary condition, do not saturate to the corresponding Page value $S_{p}$ as expected in a ergodic driven system; instead $S/S_{P}=\eta<1$ for large $n$ and large $V_{1}$ as shown in Fig. 3(a). Interestingly, $\eta$ depends on the initial Fock state; this phenomenon was interpreted to be signature of fragmentation of the Hilbert space in which case the state picked up always belong to a definite fragment (i.e a Krylov subspace) whose HSD determines $\eta$. This feature is shown in Fig. 3(b).
@@
In addition, a more experimentally accessible quantity, namely, the autocorrelation of fermion density $C_{L}(nT)=\langle\hat{n}_{L}(nT)\hat{n}_{L}(0)\rangle$ has also been studied for such systems. For ergodic system $C_{L}(nT)\to 0$ for $n\to\infty$ while for a system with HSF, it always stays above the Mazur bound $C_{b}\sim 0.125$ [133, 105]. It was found that at special frequencies and large drive amplitude, $C_{L}>C_{b}$; in contrast, at other frequencies or lower drive amplitudes, $C\to 0$ within a few drive cycles. This is shown in Fig. 3(c) for the model described by Eq. 18 using open boundary condition. Several other quantities of the driven fermion chain such as out-of-time correlators (OTOC) [80] and equal time correlation functions [79] shows features different from an ergodic driven system; these constitute a signature of prethermal HSF due to emergent approximate symmetries at special frequencies and high drive amplitude alluded to earlier. Some of these features, such as oscillations of OTOC indicating lack of information scrambling [80] and oscillations of correlators staring from a frozen ${\mathbb{Z}}_{2}$ state [79] have no analogue in fragmentation seen in equilibrium Hamiltonians.
@@
The crossover from the fragmented to ergodic behavior at the special drive frequencies can be studied as a function of the drive amplitude. The fragmented behavior persists as long as $H_{F}^{(1)}$ controls the dynamics; this sets the prethermal timescale. This timescale turns out to be exponentially sensitive to the drive amplitude at the special frequencies; thus at low drive amplitudes, the prethermal regime shrinks to a few drive cycles while at large drive cycles it can be extremely large. In between, there is a sharp crossover between the two regimes as shown in Fig. 3(d); the crossover becomes flatter with increasing system size showing the finite-sized nature of this phenomenon.
@@
Before ending this section, we note that most of the model Hamiltonians [119] studied in the context of equilibrium Hamiltonian requires imposition of additional constraints which may not be easy to implement; the prethermal version of this phenomenon does not have this requirement since the constraint is dynamically generated due to the effect of the drive. In this sense, it may have a better chance of realization in ultracold atom platforms [12, 120].
@@ III.2 Floquet scars
In classical physics, scars constitute trajectories of, for example, a moving billiard ball in a Buminovitch stadium, which do not cover the phase space completely; this represents non-ergodic behavior as known in classical dynamical systems for a long time [134]. Its implication in single-particle quantum mechanics has also been studied long back [135]. In contrast, quantum many-body scars, which are athermal near mid-spectrum eigenstates of non-integrable many-body Hamiltonians with low entanglement entropy, has been found relatively recently [16, 136]. It has been widely investigated in context Rydberg atom platforms [137, 138, 139, 140, 141, 142, 143, 144, 145], quantum spin models [146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156], interacting bosons [157, 158, 159, 160], quantum link models [161, 162, 163, 164, 165, 166], fermion models [167], open quantum systems [168, 169], and non-Hermitian models [170].
The generic conditions required for the presence of such athermal scar eigenstates in the spectrum of a many-body quantum Hamiltonian is not completely understood. In a class of these models, their presence can be understood in terms of emergence of additional symmetries [136]; this leads to a quasiparticle picture for these states. In the Rydberg atom platforms, such emergent symmetries turn out to be approximate leading to decay of scar induced oscillations of correlation functions [136, 144]. In some other cases[156, 161, 162, 163, 164, 165, 166], the presence of such scars cannot be understood from such emergent symmetries; they do not admit a simple quasiparticle picture.
To understand the presence of such emergent symmetry, one may consider a simple model, namely a spin-one XY model in magnetic field discussed in this context in Ref. 136. The Hamiltonian of this model is given by
$H=-J\sum_{\langle ij\rangle}\left(S_{i}^{+}S_{j}^{-}+{\rm h.c.}\right)-B_{0}\sum_{j}S_{j}^{z}$ (20)
@@
The key observation for obtaining scar states in this model constitutes identifying a class of excited eigenstates (bimagnons) that have similar properties. We note that the existence of such states are not guaranteed, unlike single magnon states, by charge ($S_{z}$) and total momentum conservation [136]. These states can be constructed from $|G\rangle$ by repeated application of the operator
$J_{+}$ $=$ $\sum_{\ell}(-1)^{\ell}(S_{\ell}^{+})^{2}/2$ (21)
on the ground state and are given by [136]
@@
form a spin-$L/2$ emergent ${\rm SU}(2)$ algebra. This emergent ${\rm SU}(2)$ symmetry is crucial for dynamical properties of the system. An initial state with large overlap with any of the bimagnon tower of states evolves under Hamiltonian dynamics within the bimagnon subspace due to this emergent symmetry. This feature leads to oscillatory behavior of spin-spin correlation functions; the long-time values of these correlation functions and the entanglement entropy differs from the prediction of ETH [136]. Notably, the transition between the different scar (bimagnon) eigenstates are mediated by $J^{\pm}$; these local operators implement a spectrum generating algebra $[H,J^{+}]=\omega J^{+}$ where $\omega=E_{n}-E_{n-1}=B_{0}$ and thus allows for a quasiparticle description [136].
@@
This simple picture of emergent symmetry, in its pristine form, does not hold for other, more complicated, spin models [146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156] or Rydberg atom platforms [137, 138, 139, 140, 141, 142, 143, 144, 145]. For example, the low-energy effective Hamiltonian of a Rydberg atom chain is given by [17]
$H_{\rm Ryd}$ $=$ $\sum_{j}(-\Delta\hat{n}_{j}^{z}+\Omega\sigma_{j}^{x})+V_{0}\sum_{j,j^{\prime}}\frac{\hat{n}_{j}\hat{n}_{j^{\prime}}}{|j-j^{\prime}|^{6}},$ (24)
@@
where $\hat{n}_{j}$ denotes the density operator for the Rydberg excitation on site $j$ with $\sigma_{j}^{z}=2\hat{n}_{j}-1$, $\Delta$ is the detuning parameter which can be tuned to energetically favor Rydberg excitations, $\Omega$ denotes the coupling strength between the ground and Rydberg-excited states at site $j$, and $V_{0}$ is the van der Walls repulsion strength between two excited Rydberg atoms. Such an interaction with large $V_{0}$ may exclude neighboring Rydberg excitations; this phenomenon is known as Rydberg blockade [171]. Using this property, it is possible to obtain $Z_{n}$ symmetry broken ground states of $H_{\rm Ryd}$ for large $V_{0}$ and $\Delta$ [17]. In particular, the $Z_{2}$ symmetry broken state does not allow for neighboring Rydberg excitations; in this limit, $H_{\rm ryd}$ may be described a constrained model given by [172, 173]
$H_{\rm spin}$ $=$ $\sum_{j}(-\lambda\sigma_{j}^{z}+\Omega\tilde{\sigma}_{j}^{x})$ (25)
@@
where $\tilde{\sigma}_{j}^{x}=P_{j-1}\sigma_{j}^{x}P_{j+1}$, and $P_{j}=(1-\sigma_{j}^{z})/2$ projects the spin at $j$ to $|\downarrow_{j}\rangle$. The PXP model is obtained from $H_{\rm spin}$ for $\lambda=0$ and has been shown to host quantum scar eigenstates [137, 138, 139, 140, 141, 142, 143, 144, 145]. An analysis of the model carried out in details in Refs. 144 shows that there is no exact spectrum generating algebra [144, 136]. An identification of $H_{\pm}=\sum_{j}(\tilde{\sigma}_{2j}^{\pm}+\tilde{\sigma}_{2j-1}^{\mp})$ leads to $[H_{+},H_{-}]=S_{z}+{\mathcal{O}}_{z}$ where $S_{z}$ denotes the staggered magnetization, $\sum_{j}(-1)^{j}\sigma_{j}^{z}$, and ${\mathcal{O}}_{z}$ is given by [136, 144]
${\mathcal{O}}_{z}=\sum_{j}(-1)^{j}\sigma_{j-1}^{z}\sigma_{j}^{z}\sigma_{j+1}^{z}$ (26)
@@
Figure 4: Plot of the fidelity ${\mathcal{F}}(n)=|\langle\psi(nT)|{\mathbb{Z}}_{2}\rangle|$ as a function of $n$ showing periodic persistent revivals where the dynamics is controlled by scars [top left and bottom right panels] and fast decay with no subsequent revival in their absence [bottom left panel]. An intermediate behavior indicating crossover from coherent to thermal regime is shown in the top right panel. For all plots, $\lambda_{0}=15\Omega$ and all figures are obtained from Ref. 84.
The Floquet version of this phenomenon has been studied in several works in the context of Rydberg atom platforms [84, 85, 86, 87, 88], fractonic models [83], and trimerized Kagome lattice models [89]. For Rydberg atom platform, the drive was implemented by making the longitudinal magnetic field term $H_{0}=-\lambda\sigma_{j}^{z}$ in $H_{\rm spin}$ (Eq. 25) time dependent; $\lambda$ was varied a function of time using a square-pulse [84, 86, 85], periodic kicks [87] or continuous cosine drive protocol [88]. In each of these cases, scar eigenstates were found in the Floquet quasienergy spectrum of the driven model; they were found to influence the dynamics of the model and leave their imprint on correlation functions and entanglement entropy.
For the square pulse protocol, it was found in Refs. 84, 85, that the Floquet Hamiltonian admits scar states which has large overlap $|{\mathbb{Z}}_{2}\rangle$ at high and intermediate frequency range except at special drive frequencies where the first order Floquet Hamiltonian vanishes. For the square-pulse protocol the first order Floquet Hamiltonian was found to be of the PXP form; its perturbative analytical expression in the large drive amplitude regime is given by [84]
@@
where $\lambda_{0}$ is the drive amplitude and $T$ is the time period of the drive. $H_{F}^{(1)}$ vanishes at special drive frequencies, $\omega_{n}^{\ast}$ for which $\lambda_{0}T=4n\pi$ ($n\in Z$); away from these frequencies, at high drive amplitude, it controls the dynamics up to a large prethermal time scale. Since $H_{F}^{(1)}$ has the same form as $H_{\rm PXP}$, albeit with a rescaled $\Omega$ and an unimportant rotation in spin space, it’s eigenspectrum hosts quantum scars. These scars lead to oscillatory dynamics of correlation functions, similar to the static case; such a behavior can be understood in context of emergent approximate $SU(2)$ symmetry as discussed above. Note that in the Floquet context this emergent symmetry is approximate both due to non-closure of the ${\rm SU}(2)$ algebra and also from the fact that the higher order terms in the Floquet Hamiltonian do not respect such symmetry. The latter effect becomes particularly apparent at the special frequencies where $H_{F}^{(1)}$ vanishes; for these frequencies one sees fast, ETH predicted, thermalization as can be seen in Fig. 4. This allows for possibility of controlling quantum coherence using the drive frequency as a tuning parameter.
Figure 5: (a) Plot of $S_{L/2}$ for the eigenstates of $H_{F}$ for $L=26$ and $\hbar\omega_{D}/\Omega=8.5$ at $\lambda_{0}/\Omega=15$. The eigenstates with overlap $>0.01$ with $|0\rangle(|\mathbb{Z}_{2}\rangle)$ are shown using red (green) circles. These states are distinct and coexist at this drive frequency. (b) Plot of $O_{22}=\langle\hat{n}_{2}\hat{n}_{4}\rangle$ as a function of Floquet eigenstate quasienergies $E_{F}$ (in units of $\Omega$). The violet dashed line indicates the ETH predicted value of $O_{22}$ at a temperature $T_{0}=E_{F}/k_{B}$. These figures, adapted from Ref. 86, establishes coexistence of two different type of scar states in the driven PXP chain.
The Floquet scars found in Refs. 84, 85 is continually connected to the static scars of the PXP model in the high drive frequency limit. However, another class of scars which do not have any equilibrium analogue has also been found at intermediate frequencies [86]. Unlike their counterparts which have large overlap with $|{\mathbb{Z}}_{2}\rangle$ Fock states, these show significant overlap with $|0\rangle$ state. They coexists with the ${\mathbb{Z}}_{2}$ scars at intermediate drive frequency range $8\leq\hbar\omega_{D}/\Omega\leq 10$. Such scars necessarily arise from the combined effect of higher order terms in the Floquet Hamiltonian; the first of such terms, derived using FPT in the high drive amplitude regime, is given by [86]
@@
It turns out that eigenstates of $H_{F}^{(1)}+H_{F}^{(3)}$ hosts the additional scars which has large overlap with $|0\rangle$; these therefore do not have any equilibrium analogue. Their coexistence with the ${\mathbb{Z}}_{2}$ scar state is shown in Fig. 5; the corresponding oscillatory dynamics of correlation function $\langle\hat{n}_{j}\hat{n}_{j+2}\rangle$ starting from the $|0\rangle$ state due to such scars has been shown in Ref. 86. The existence of these Floquet scars does not seem to be attributable to any emergent symmetry of the Floquet Hamiltonian.
More recently, a period-kick protocol has been studied in this context [87]. It has been shown that such a drive can induce, at specific value of the kick amplitude and frequency, a subharmonic response which can be interpreted as a signature of prethermal $Z_{2}$ time crystal. The difference of this phenomenon with its more standard counterpart (see Sec. IV for a detailed discussion) is that such a response occurs only when the starting state is a Neel state and the oscillations, which occur between the Neel state and its time-reversed partner, are controlled by $Z_{2}$ scars. These scar states share the same approximate emergent symmetry discussed earlier. In contrast, scar induced oscillations for both the Neel and the vacuum initial states has also been found for a cosine drive protocol in the non-perturbative regime [88]. As in Ref. 86, the origin of these scars can not be understood from an emergent symmetry principle.

### v2

@@ III.1 Prethermal Hilbert-space fragmentation
For $\gamma_{1}=n\pi$, or $V_{1}=2n\hbar\omega_{n}^{\ast}$, $\alpha(\gamma_{1})=0$. At these special frequencies, $H_{F}^{(1)}$ represents a model of interacting 1D spinless fermions with constrained hopping; such a model, in equilibrium, leads to strong Hilbert-space fragmentation hsf4 . However, higher order terms in $H_{F}$ (Ref. sg2 ) do not respect such constraints. Most importantly, the presence of such constraints leads to two approximate emergent conservations; $[H_{F}^{(1)},n_{d}]=0$ and $[H_{F}^{(1)},n_{d}^{s}]=0$, where $n_{d}=\sum_{j}\hat{n}_{j}\hat{n}_{j+1}$ and $n_{d}^{s}=\sum_{j}(-1)^{j}\hat{n}_{j}\hat{n}_{j+1}$. Note that these conservation laws are approximate as higher order terms in $H_{F}$ do not respect them. These approximate emergent conservation laws are crucial in obtaining a prethermal Hilbert-space fragmentation phase at special frequencies.
@@
The signature for such Hilbert-space fragmentation becomes apparent in several quantities. For example, one expects the entanglement entropy for a generic initial state of an ergodic driven system to eventually saturate to its Page value $S_{p}$. This can be thought as the average entropy for a generic ergodic many-body system; it is also the typical entropy for large enough systems and depends on the HSD of the system (see Refs. pageref, ; dona1, for details). However, at these special frequencies and for high drive amplitude, the half-chain entanglement entropy $S(nT)$, of an initial Fock state after $n$ drive cycles does not saturate to the corresponding Page value $S_{p}$ as expected in a ergodic driven system. Instead, $S/S_{P}=\eta<1$ for large $n$ and large $V_{1}$ as shown in the top left panel of Fig. 3. Interestingly, $\eta$ depends on the initial Fock state. This phenomenon was interpreted to be signature of fragmentation of the Hilbert space in the computational Fock basis in which case the state picked up always belongs to a definite fragment (i.e a Krylov subspace) whose HSD determines $\eta$. This feature is shown in the top right panel of Fig. 3.
@@
In addition, a more experimentally accessible quantity, namely, the autocorrelation of fermion density $C_{L}(nT)=\langle\hat{n}_{L}(nT)\hat{n}_{L}(0)\rangle$ has also been studied for such systems. For ergodic system $C_{L}(nT)\to 0$ for $n\to\infty$ while for a system with Hilbert-space fragmentation, it always stays above the Mazur bound $C_{b}\sim 0.125$ majur1 ; hsf5 . It was found that at special frequencies and large drive amplitude, $C_{L}>C_{b}$. In contrast, at other frequencies or lower drive amplitudes, $C\to 0$ within a few drive cycles. This is shown in the bottom left panel Fig. 3 for the model described by Eq. 18 using open boundary conditions.
Several other quantities of the driven fermion chain such as out-of-time correlators (OTOC) sg2 and equal time correlation functions sg1 show features different from an ergodic driven system. These constitute a signature of prethermal Hilbert-space fragmentation due to emergent approximate symmetries at special frequencies and high drive amplitude alluded to earlier. Some of these features, such as oscillations of OTOC indicating lack of information scrambling sg2 and oscillations of correlators staring from a frozen ${\mathbb{Z}}_{2}$ state sg1 have no analogue in fragmentation seen in equilibrium Hamiltonians.
@@
The crossover from the fragmented to ergodic behavior at the special drive frequencies can be studied as a function of the drive amplitude. The fragmented behavior persists as long as $H_{F}^{(1)}$ controls the dynamics; this sets the prethermal timescale. This timescale turns out to be exponentially sensitive to the drive amplitude at the special frequencies. Thus at low drive amplitudes, the prethermal regime shrinks to a few drive cycles while at large drive cycles it can be extremely large. In between, there is a sharp crossover between the two regimes as shown in the bottom right panel of Fig. 3; the crossover becomes flatter with increasing system size showing the finite-sized nature of this phenomenon.
@@
Before ending this section, we note that most of the model Hamiltonians hsfrev , studied in the context of equilibrium Hamiltonian, require imposition of additional constraints which may not be easy to implement; the prethermal version of this phenomenon does not have this requirement since the constraint is dynamically generated due to the effect of the drive. In this sense, it may have a better chance of realization in ultracold atom platforms rev12 ; hsfexp .
@@ III.2 Floquet scars
In classical physics, scars constitute trajectories of, for example, a moving billiard ball in a Buminovitch stadium, which do not cover the phase space completely. This represents non-ergodic behavior as known in classical dynamical systems for a long time scarcl . Its implication in single-particle quantum mechanics has also been studied long back scarqm1 . In contrast quantum many-body scars are athermal, near mid-spectrum, eigenstates of non-integrable many-body Hamiltonians with low entanglement entropy. They have been discovered relatively recently exp3 ; scarrev . Since then, quantum many-body scars have been widely investigated in context Rydberg atom platforms scarpxp1 ; scarpxp2 ; scarpxp3 ; scarpxp4 ; scarpxp5 ; scarpxp6 ; scarpxp7 ; scarpxp8 ; scarpxp9 , quantum spin models scarspin1 ; scarspin2 ; scarspin3 ; scarspin4 ; scarspin5 ; scarspin6 ; scarspin7 ; scarspin8 ; scarspin9 ; scarspin10 ; scarspin11 ; scarspin12 , interacting bosons scarbose1 ; scarbose2 ; scarbose3 ; scarbose4 ; scarbose5 , quantum link models scarlink1 ; scarlink2 ; scarlink3 ; scarlink4 ; scarlink5 ; scarlink6 ; scarlink7, fermion models scarferm1 , open quantum systems scaropen1 ; scaropen2 , non-Hermitian models scarnh1 , and random network models mar1 .
The generic conditions required for the presence of such athermal scar eigenstates in the spectrum of a many-body quantum Hamiltonian is not completely understood. In a class of these models, their presence can be understood in terms of emergence of additional symmetries scarrev ; these lead to a quasiparticle picture for these states. In the Rydberg atom platforms, such emergent symmetries turn out to be approximate leading to decay of scar induced oscillations of correlation functions scarrev ; scarpxp8 . In some other casesscarspin11 ; scarlink1 ; scarlink2 ; scarlink3 ; scarlink4 ; scarlink5 ; scarlink6 , the presence of such scars cannot be understood from such emergent symmetries; they do not admit a simple quasiparticle picture.
To understand the presence of such emergent symmetry, one may consider a simple model, namely a spin-one XY model in magnetic field discussed in this context in Ref. scarrev, . The Hamiltonian of this model is given by
$H=-J\sum_{\langle ij\rangle}\left(S_{i}^{+}S_{j}^{-}+{\rm h.c.}\right)-B_{0}\sum_{j}S_{j}^{z},$ (20)
@@
The key observation for obtaining scar states in this model constitutes identifying a special class of excited eigenstates (bimagnons) that have similar properties. We note that the existence of such states is not guaranteed, unlike single magnon states, by charge ($S_{z}$) and total momentum conservation scarrev . These states can be constructed from $|G\rangle$ by repeated application of the operator
$J_{+}$ $=$ $\sum_{\ell}(-1)^{\ell}(S_{\ell}^{+})^{2}/2,$ (21)
on the ground state and are given by scarrev
@@
This emergent ${\rm SU}(2)$ symmetry is crucial for dynamical properties of the system. An initial state with large overlap with any of the bimagnon tower of states evolves under Hamiltonian dynamics within the bimagnon subspace due to this emergent symmetry. This feature leads to oscillatory behavior of spin-spin correlation functions; the long-time values of these correlation functions and the entanglement entropy differs from the prediction of ETH scarrev . Notably, the transition between the different scar (bimagnon) eigenstates are mediated by $J^{\pm}$. These local operators implement a spectrum generating algebra $[H,J^{+}]=\omega J^{+}$ where $\omega=E_{n}-E_{n-1}=B_{0}$ and thus allows for a quasiparticle description scarrev .
@@
This simple picture of emergent symmetry, in its pristine form, does not hold for other, more complicated, spin models scarspin1 ; scarspin2 ; scarspin3 ; scarspin4 ; scarspin5 ; scarspin6 ; scarspin7 ; scarspin8 ; scarspin9 ; scarspin10 ; scarspin11 or Rydberg atom platforms scarpxp1 ; scarpxp2 ; scarpxp3 ; scarpxp4 ; scarpxp5 ; scarpxp6 ; scarpxp7 ; scarpxp8 ; scarpxp9 ; scarspin12 . For example, the low-energy effective Hamiltonian of a Rydberg atom chain is given by exp4
$H_{\rm Ryd}$ $=$ $\sum_{j}(-\Delta\hat{n}_{j}^{z}+\Omega\sigma_{j}^{x})+V_{0}\sum_{j,j^{\prime}}\frac{\hat{n}_{j}\hat{n}_{j^{\prime}}}{|j-j^{\prime}|^{6}},$
@@
where $\hat{n}_{j}$ denotes the density operator for the Rydberg excitation on site $j$ with $\sigma_{j}^{z}=2\hat{n}_{j}-1$. Here $\Delta$ is the detuning parameter which can be tuned to energetically favor Rydberg excitations, $\Omega$ denotes the coupling strength between the ground and Rydberg-excited states at site $j$, and $V_{0}$ is the van der Walls repulsion strength between two excited Rydberg atoms.
Such an interaction with large $V_{0}$ may exclude neighboring Rydberg excitations; this phenomenon is known as Rydberg blockade blref1 . Using this property, it is possible to obtain $Z_{n}$ symmetry broken ground states of $H_{\rm Ryd}$ for large $V_{0}$ and $\Delta$ exp4 . In particular, the $Z_{2}$ symmetry broken state does not allow for neighboring Rydberg excitations; in this limit, $H_{\rm ryd}$ may be described a constrained model given by ks1 ; ks2
$H_{\rm spin}$ $=$ $\sum_{j}(-\lambda\sigma_{j}^{z}+\Omega\tilde{\sigma}_{j}^{x}),$ (25)
where $\tilde{\sigma}_{j}^{x}=P_{j-1}\sigma_{j}^{x}P_{j+1}$, and $P_{j}=(1-\sigma_{j}^{z})/2$ projects the spin at $j$ to $|\downarrow_{j}\rangle$. The PXP model is obtained from $H_{\rm spin}$ for $\lambda=0$.
@@
The PXP model has been shown to host quantum scar eigenstates scarpxp1 ; scarpxp2 ; scarpxp3 ; scarpxp4 ; scarpxp5 ; scarpxp6 ; scarpxp7 ; scarpxp8 ; scarpxp9 . An analysis of the model carried out in details in Refs. scarpxp8, shows that there is no exact spectrum generating algebra scarpxp8 ; scarrev . An identification of $H_{\pm}=\sum_{j}(\tilde{\sigma}_{2j}^{\pm}+\tilde{\sigma}_{2j-1}^{\mp})$ leads to $[H_{+},H_{-}]=S_{z}+{\mathcal{O}}_{z}$ where $S_{z}$ denotes the staggered magnetization, $\sum_{j}(-1)^{j}\sigma_{j}^{z}$, and ${\mathcal{O}}_{z}$ is given by scarrev ; scarpxp8
${\mathcal{O}}_{z}=\sum_{j}(-1)^{j}\sigma_{j-1}^{z}\sigma_{j}^{z}\sigma_{j+1}^{z}.$ (26)
The presence of non-zero ${\mathcal{O}}_{z}$ indicates that the emergent ${\rm SU}(2)$ algebra is not closed; it constitutes the realization of an approximate ${\rm SU}(2)$ algebra scarpxp8 .
@@
Figure 4: Plot of the fidelity ${\mathcal{F}}(n)=|\langle\psi(nT)|{\mathbb{Z}}_{2}\rangle|$ as a function of $n$ showing periodic persistent revivals where the dynamics is controlled by scars [top left and bottom right panels] and fast decay with no subsequent revival in their absence [bottom left panel]. An intermediate behavior indicating crossover from coherent to thermal regime is shown in the top right panel. For all plots, $\lambda_{0}=15\Omega$ and all figures are obtained from Ref. bm1, .
The Floquet version of this phenomenon has been studied in several works in the context of Rydberg atom platforms bm1 ; mituza1 ; bm2 ; lukinsc1 ; papic1 , fractonic models pretko1 , and trimerized Kagome lattice models liu1 . For Rydberg atom platform, the drive was implemented by making the longitudinal magnetic field term $H_{0}=-\lambda\sigma_{j}^{z}$ in $H_{\rm spin}$ (Eq. 25) time dependent; $\lambda$ was varied a function of time using a square-pulse bm1 ; bm2 ; mituza1 , periodic kicks lukinsc1 or continuous cosine drive protocol papic1 . In each of these cases, scar eigenstates were found in the Floquet quasienergy spectrum of the driven model; they were found to influence the dynamics of the model and leave their imprint on correlation functions and entanglement entropy.
For the square pulse protocol, it was found in Refs. bm1, ; mituza1, , that the Floquet Hamiltonian admits scar states that have large overlap $|{\mathbb{Z}}_{2}\rangle$ at high and intermediate frequency range except at special drive frequencies where the first order Floquet Hamiltonian vanishes. For the square-pulse protocol the first order Floquet Hamiltonian was found to be of the PXP form. It’s perturbative analytical expression in the large drive amplitude regime is given by bm1
@@
where $\lambda_{0}$ is the drive amplitude and $T$ is the time period of the drive. $H_{F}^{(1)}$ vanishes at special drive frequencies, $\omega_{n}^{\ast}$ for which $\lambda_{0}T=4n\pi$ ($n\in Z$); away from these frequencies, at high drive amplitude, it controls the dynamics up to a large prethermal time scale.
Since $H_{F}^{(1)}$ has the same form as $H_{\rm PXP}$, albeit with a rescaled $\Omega$ and an unimportant rotation in spin space, it’s eigenspectrum hosts quantum scars. These scars lead to oscillatory dynamics of correlation functions, similar to the static case; such a behavior can be understood in the context of an emergent approximate $SU(2)$ symmetry as discussed above. Note that in the Floquet context this emergent symmetry is approximate both due to the non-closure of the ${\rm SU}(2)$ algebra and also from the fact that the higher order terms in the Floquet Hamiltonian do not respect such symmetry. The latter effect becomes particularly apparent at the special frequencies where $H_{F}^{(1)}$ vanishes; for these frequencies one sees fast, ETH predicted, thermalization as can be seen in Fig. 4. This allows for possibility of controlling quantum coherence using the drive frequency as a tuning parameter.
Figure 5: (a) Plot of $S_{L/2}$ for the eigenstates of $H_{F}$ for $L=26$ and $\hbar\omega_{D}/\Omega=8.5$ at $\lambda_{0}/\Omega=15$. The eigenstates with overlap $>0.01$ with $|0\rangle(|\mathbb{Z}_{2}\rangle)$ are shown using red (green) circles. These states are distinct and coexist at this drive frequency. (b) Plot of $O_{22}=\langle\hat{n}_{2}\hat{n}_{4}\rangle$ as a function of Floquet eigenstate quasienergies $E_{F}$ (in units of $\Omega$). The violet dashed line indicates the ETH predicted value of $O_{22}$ at a temperature $T_{0}=E_{F}/k_{B}$. These figures, adapted from Ref. bm2, , establishes coexistence of two different type of scar states in the driven PXP chain.
The Floquet scars found in Refs. bm1, ; mituza1, are continually connected to the static scars of the PXP model in the high drive frequency limit. However, another class of scars which do not have any equilibrium analogue has also been found at intermediate frequencies bm2 . Unlike their counterparts which have large overlap with $|{\mathbb{Z}}_{2}\rangle$ Fock states, these show significant overlap with $|0\rangle$ state. They coexists with the ${\mathbb{Z}}_{2}$ scars at intermediate drive frequency range $8\leq\hbar\omega_{D}/\Omega\leq 10$. Such scars necessarily arise from the combined effect of higher order terms in the Floquet Hamiltonian. The first of such terms, derived using FPT in the high drive amplitude regime, is given by bm2
@@
It turns out that the eigenstates of $H_{F}^{(1)}+H_{F}^{(3)}$ host additional scars which has large overlap with $|0\rangle$; these therefore do not have any equilibrium analogue. Their coexistence with the ${\mathbb{Z}}_{2}$ scar state is shown in Fig. 5; the corresponding oscillatory dynamics of correlation function $\langle\hat{n}_{j}\hat{n}_{j+2}\rangle$ starting from the $|0\rangle$ state due to such scars has been shown in Ref. bm2, . The existence of these Floquet scars does not seem to be attributable to any emergent symmetry of the Floquet Hamiltonian.
More recently, a period-kick protocol has been studied in this context lukinsc1 . It has been shown that such a drive can induce, at specific value of the kick amplitude and frequency, a subharmonic response which can be interpreted as a signature of prethermal $Z_{2}$ time crystal. The difference of this phenomenon with its more standard counterpart (see Sec. IV for a detailed discussion) is that such a response occurs only when the starting state is a Neel state. The corresponding oscillations, which occur between the Neel state and its time-reversed partner, are controlled by $Z_{2}$ scars. These scar states share the same approximate emergent symmetry discussed earlier. In contrast, scar induced oscillations for both the Neel and the vacuum initial states have also been found for a cosine drive protocol in the non-perturbative regime papic1 . As in Ref. bm2, , the origin of these scars can not be understood from an emergent symmetry principle.